### v1

@@ 6 $\beta$ Cep ensemble asteroseismology
To gain deeper insights into the population of the Gaia-observed $\beta$ Cep stars, we used the identification of the dominant modes. In the following, we characterise the ensemble and subject its stars to grid modelling.
@@ 6.1 Distributions of the identified modes
The mean behaviour of the identified modes of the full sample over the wavelength is very similar to the hand-selected sample in Fig. 8. However, we find a slightly clearer separation between the two considered non-radial modes. Notably the distributions of $A_{T}/A_{\mathrm{BP}}$ for $l=1,2$ shown in Fig. 19 are bimodal in the TESS amplitude but not in the other photometric bands.
Figure 9: Amplitude ratio diagram based on $A_{\mathrm{RP}}/A_{\mathrm{BP}}$ and $A_{T}/A_{\mathrm{BP}}$ similar to Fig. 6. The colour-code shows the mode degree as determined with our probabilistic framework: $l=0$ (green), $l=1$ (blue), $l=2$ (orange). Grey symbols denote stars with $p<0.5$. The symbol size is proportional to the probability of the mode identification. Stars marked with black outlines constitute the underlying sample from Fig. 8.
@@
When going into details, we find a complicated pattern with several overlapping regions. Due to the three dimensionality of the distributions these overlap regions create fuzzy borders in the two dimensional projection in Fig. 9. Stars to which our algorithm assigned low probabilities in the mode identification belong to the overlapping regions between the different degrees. An interesting feature is the concentration of $l=2$ modes near $A_{\mathrm{RP}}/A_{\mathrm{BP}}\approx 1$ and $A_{\mathrm{T}}/A_{\mathrm{BP}}\approx 0.5$ which is clearly separated from the main population at the top. Similarly, the group of high probability $l=1$ modes at the bottom of Fig. 9 is disjoint from the main $l=1$ population. These two regions create the bimodalities in the $A_{\mathrm{T}}/A_{\mathrm{BP}}$ amplitude ratios in Fig. 19. From our TESS data, we find for some of these stars obvious multiplets in which the Gaia observed frequency (and often main TESS frequency) is not the central frequency. Hence, the multi-colour photometry also carries information on the azimuthal order $m$.
When placing the stars with identified modes in the HRD, we find the $\beta$ Cep stars in which the radial mode is the dominant mode to be mainly of lower mass ($\lesssim 15\,M_{\sun}$) while non-radial pulsators can be observed throughout the whole $\beta$ Cep mass range (cf. Fig. 20). Two stars with radial modes stand out due to their relatively high luminosity and mass. Notably, the radial pulsators seem to concentrate in the later half of the main sequence in the overlap region between the $\beta$ Cep and SPB instability strips.
@@ 6.2 Rotationally split multiplets and asymmetries
As introduced in Sect. 5.2, rotation can split non-radial modes into multiplets with splittings proportional to the envelope rotation rate. In Fig. 10, we show the observed splittings against the estimated surface rotation rate. As expected, a clear positive correlation between the rotation rate and the splitting is observed. Yet, we recall that $\beta$ Cep stars can have strong radial differential rotation (Aerts et al., 2003; Pamyatnykh et al., 2004; Burssens et al., 2023) and the surface rotation rate may differ from the internal rotation rate at the fractional radius where the mode is most sensitive.
@@
Figure 10: Observed mode splittings against the projected surface rotation rate ($f_{\mathrm{rot}}\sin i$, thin outlines) or photometric surface rotation rate ($f_{\mathrm{rot}}$, solid outlines). The three lines show the widths of splittings equal to the surface rotation rate (solid) as well as half (dotted) and twice (dashed) the surface rotation rate. We note that the internal rotation rate at the position of maximum probing power might be different from the surface rotation. The uncertainties on the splittings indicate the maximum difference of the individual splittings to the mean splitting of the multiplet and are a measure for the asymmetry. The colours show our manual classification into $l=1$ (blue) and $l=2$ (orange) modes.
@@
We also mark the absolute asymmetries (maximum difference to the mean splitting) in Fig. 10. For most of the stars this asymmetry is small, though it is pronounced for a minority. A purely rotationally split multiplet would consist of evenly and symmetrically spaced components if the Coriolis force, due to the mean envelope rotation, has only a small influence, allowing for a first-order pertubative approach. However, second order-terms are relevant for $\beta$ Cep stars and include rotational deformation of the star (Saio, 1981). This effect shifts the split frequencies asymmetrically with respect to the central frequency (Dziembowski & Goode, 1992; Goupil et al., 2000). In addition, properties such as magnetic fields (e.g. Bugnet et al., 2021) or binarity (e.g. Guo et al., 2024) also influence the asymmetry of split multiplets, leading to asymmetries in the observed frequency spectra and enabling detailed asteroseismic inferences (e.g. Vanlaer et al., in prep.).
@@
Figure 11: Asymmetries of 48 complete rotationally split multiplets $\delta_{A}$ against their central frequency $f_{0}$ colour-coded by the projected surface rotation rate. Square symbols mark stars with measured photometric surface rotation rates ($f_{\mathrm{rot}}$). Symbols connected by a line belong to the same ($l=2$) multiplet. The measurement uncertainties are within the symbol sizes.
@@
with $f_{0}$ the central frequency and $f_{\pm m}$ the frequencies of azimuthal order $m=\pm 1$. In Fig. 11, we show the asymmetries for 48 stars in our sample against the central frequency $f_{0}$ of the multiplet. Most of the splittings are positive as expected for a simple rotational deformation (Guo et al., 2024). The splittings shown in Fig. 11 are only the splittings most obvious and close to the main oscillation mode. Several stars show additional splittings which will be analysed elsewhere.
Negative asymmetries in $\beta$ Cep stars hint towards internal magnetic fields. Our understanding of core magnetism in massive stars is still limited (Lecoanet et al. 2022; Vanlaer et al., in prep.), despite its potential impact on stellar evolution. A large sample of measured internal magnetic fields from asteroseismology would deliver constraints for evolutionary models, similarly to the recent advances in red giants (Li et al., 2023).
@@ 6.3 Modelling the population of $\beta$ Cep stars
To understand the population of $\beta$ Cep stars better and gain insights into the distributions of their physical properties with respect to the selected main pulsation mode, we applied grid modelling. We used the $\beta$ Cep MESA stellar evolution grid presented by Burssens et al. (2023). The grid was calculated for masses between 9 $M_{\sun}$ and 21.5 $M_{\sun}$ at solar metallicity with envelope mixing by internal gravity waves set to $D=1000\,$cm2 s-1 at the base of the envelope and increasing outwards as $\rho^{-1}$ and an exponentially decaying diffusive core overshoot with the parameter range $0.005\leq f_{\mathrm{ov}}\leq 0.035$. The $\beta$ Cep grid mimics the effect of rotational mixing by including core overshoot and envelope mixing. For details of the physical implementation, we refer the reader to Burssens et al. (2023).
@@
The required mode identification for the modelling is not limited to the degree $l$ but we require the radial order $n$ and azimuthal order $m$ as well. All models will be fit in the azimuthal order $m=0$ necessitating us to identify the central peak of split modes. For complete multiplets the choice is obvious. In incomplete multiplets, we assumed in most cases the highest peak to be the at $m=0$. Exceptions from this rule were for example made when the highest amplitude TESS frequency does not agree with the Gaia frequency.
@@
The radial order can neither be deduced from the observations nor from a direct fit of the model. Rather, we fitted model with the radial orders $n\in\{0,1,2\}$ for radial pulsators and $n\in\{-2,-1,0,1,2\}$ for non-radial modes [^4] The f-mode ($n=0$) only exists for $l=2$ as the frequency goes to zero for dipole modes (e.g. Takata, 2012). using a $\chi^{2}$ minimisation over $\log T_{\mathrm{eff}}$, $\log L$, and $f_{0}$. In principle, the radial order can be larger but higher order modes have not been securely established in $\beta$ Cep stars, hence we did not probe them.
@@
Stars without a model within the $1\sigma$ uncertainty ellipse (for which $\Delta f_{0}$ was inflated to 0.5 d-1, see below) were excluded from the subsequent analysis, reducing the number of modelled stars to 164. For each star, we chose the best fitting model for each $n$ and compared the position of its modes to the observed frequency spectra. Our focus laid on the relative positions of the main pulsation mode and other visible modes with respect to the fitted model. In most but not all cases our choice of $n$ coincides with the lowest $\chi^{2}$ value among the set of best fitting models. We show one illustrative example for the identification in Appendix B, including a detailed reasoning for our choice of $n$.
@@
We find the majority of the radial pulsators to pulsate in their fundamental mode (27 out of 36 modelled radial pulsators). Seven radial pulsators pulsate in their first overtone and two stars are likely second overtone pulsators. A similar picture emerges for the 128 non-radial pulsators. The dominant pulsation mode in 72 stars is the $p_{1}$ mode ($n=1$) and 28 pulsate in $g_{1}$ ($n=-1$). Only 21 stars pulsate in their second-order mode with 18 identified $p_{2}$ and three $g_{2}$ pulsators. The remaining seven stars are likely f-mode pulsators.
@@
For the final model selection and uncertainty estimation, we employed a MCMC approach with a $\chi^{2}$ merit function, similar to our work in Fritzewski et al. (2024), to which we refer for details. Since we rely on a grid of stellar models with fixed input physics on some aspects (notably the level of envelope mixing, where the low-order modes have dominant probing power), we allow for an overall mismatch between observed and fitted frequencies (see Aerts et al., 2018, for an extensive discussion in the case of g modes). In order to encapsulate this theoretical uncertainty caused by the fixed yet uncertain input physics adopted in the model grid, we include an artificially inflated frequency uncertainty of 0.5 d-1. This procedure allows us to recognise correspondence between the observed and predicted frequency pattern of the low-order modes and thus encapsulates any systematic frequency shifts between the model predictions and the observed frequency $f_{0}$ (cf.  the procedure explained in in Appendix B).
Figure 12: Distribution of $\beta$ Cep parameters for the modelled stars. Each panel contains the histograms for the full sample (grey) and degree $l=0$ (green), $l=1$ (blue), and $l=2$ (orange). The histograms are normalised to the number of stars within the group. The mass distribution in the top panel shows that the majority of stars are found in the lower-mass end of the instability strip. The second panel gives the distribution of the normalised core hydrogen fraction $X_{c}^{\prime}$. We find the majority of stars in the second half of their main sequence lifetime. We note the inverted x-axis representing evolutionary stage. The third panel shows the actual stellar age distribution. The fractional mass of the convective core in the fourth panel shows a concentration for lower core masses in agreement with the top panel and the HRDs in this work. The radius distribution in the penultimate panel is peaked at intermediate radii. The $K$-values in the bottom panel show the dipole and quadrupole modes to have dominant radial over tangential displacements at the stellar surface.
@@
The histograms in Fig. 12 show distributions of the obtained masses, normalised core hydrogen fraction ($X_{c}^{\prime}$), age, relative convective core mass, radius, and horizontal versus radial displacement of the dominant mode. As expected from the distribution of $\beta$ Cep stars in the HRD in Fig. 3, the number of $\beta$ Cep stars declines with mass, with most $\beta$ Cep stars found in the range $9\leq M_{\mathrm{ini}}/M_{\sun}\leq 13$. In agreement with Fig. 20, radial pulsators are mostly found among the lower mass $\beta$ Cep stars.
The distribution of $X_{c}^{\prime}$ shows more structure in particular when focusing on the pulsation degree. Radial pulsators are constrained to the central region of the probed core hydrogen fraction, while non-radial pulsators can be found over the whole range. Notably the quadrupole pulsators follow a bimodal distribution with a deficiency near $X_{c}^{\prime}=0.5$. This may be connected to the phenomenon of avoided crossings for such modes in the evolved stages, in line with the findings by Mazumdar et al. (2006) for the evolved $\beta\,$Cep star $\beta\,$CMa. The stellar ages do not follow the same distribution but are similarly distributed for all three pulsation degrees.
@@
We do not show a distribution of the core overshoot because most of the stars are modelled with $f_{\mathrm{ov}}\approx 0.02$, which corresponds to the central value in the grid. Yet, the uncertainties span the whole range of $f_{\mathrm{ov}}$ covered by the grid, which can be explained by the limited sensitivity of the observed modes to the near-core region and the occurrence of degeneracies between the mass, metallicity, and core overshoot when relying on just one mode (e.g., Ausseloos et al., 2004). The estimated core overshoot values result in a value for the mass-normalised convective mass ($M_{\mathrm{core}}/M_{\star}$). This inferred parameter is the more interesting quantity from a stellar evolution point of view and it is a function of both the luminosity and age. The majority of stars has a value in the range $0.2\leq M_{\mathrm{core}}/M_{\star}\leq 0.3$. The few stars with a relatively high convective core mass belong to the higher-mass young O-stars well separated in the HRD (cf. Fig. 3).
Figure 13: Normalised convective core mass $M_{\mathrm{core}}/M_{\mathrm{ini}}$ over the normalised core hydrogen mass fraction $X_{c}^{\prime}$. The stars are colour-coded by their primary degree: $l=0$ (green), $l=1$ (blue), and $l=2$ (orange). The shape depends on the radial order. The g modes are indicated by diamonds. We use thin diamonds for g2 modes and wider ones for g1 modes. Radial fundamental modes are shown with triangles; $n=1$ of both radial and non-radial nature are indicated by circles and $n=2$ modes by squares. The grey tracks in the background show the evolution for the indicated masses with $f_{\mathrm{ov}}=0.005$ (solid, labelled at low $X_{c}^{\prime}$) and $f_{\mathrm{ov}}=0.035$ (dashed, labelled at high $X_{c}^{\prime}$). We note the inverted x-axis to represent the evolutionary stage. The typical uncertainty is indicated in the upper right corner.
@@
TIC GaiaDR3 source ID $X_{c}^{\prime}$ $M_{ini}$ $\log t$ $M_{\mathrm{core}}/M$
@@
As a final distribution, we show in the bottom panel the distribution for the so-called $K-$value, defined as $K\equiv GM/(2\pi f_{0})^{2}R^{3}$, for the non-radial modes. This dimensionless quantity is connected with the boundary conditions for the pulsation equations (Aerts et al., 2010, Chapter 3). In the case of non-radial modes, it represents the ratio between the horizontal and radial displacement of the fluid elements at the stellar surface caused by the mode (as this ratio is zero for radial modes, these are not included in the bottom panel of the figure). This parameter has high values above unity (typically between 10 and 100) for high-order g modes of observed SPB stars (De Cat & Aerts, 2002), as their modes dominantly cause horizontal displacements. Here, as expected for low-order modes, we find the opposite with dominantly radial displacements but with non-negligible contributions from horizontal motions. The values for $K$ range from about 0.03 to about 0.2 for the dominant mode of most of the Gaia $\beta$ Cep stars in our sample.
To unravel the correlation between the convective core mass and the core hydrogen fraction, we show these two parameters in Fig. 13. The overall distribution shows clearly the distinction between stars with $M_{\mathrm{core}}/M_{\star}\leq 0.3$ and these with more massive convective cores. The stars with higher relative core masses are typically more massive and younger as expected from stellar evolution theory. As seen from the evolutionary tracks of different masses, only the lowest mass $\beta$ Cep stars can be found among full breadth of the main sequence evolution, a consequence of the position of the instability strip lying closer to the zero-age main sequence for these stars. Notably, the bimodality in $X_{c}^{\prime}$ for quadrupole pulsators is present for all masses but most pronounced in the lower mass range. In particular the younger quadrupole pulsators are low-order g-mode pulsators, while their older counterparts are a mix of g- and p-mode pulsators.
In summary, we successfully modelled 164 $\beta$ Cep stars with a grid modelling approach and provide distributions of their masses, ages, convective core masses, radii. As such, these distribution provide a first insight into the properties of the population of $\beta$ Cep stars. Of particular interest for future investigations on the mode excitation of this population is the connection between the stellar age and the degree of the dominant mode, if any. We give the best fitting model properties in Table 2. A discussion of the accuracy of the asteroseismic parameters in comparison with the input values from Gaia can be found in Appendix C.1.

### v2

@@ 6 $\beta$ Cep ensemble asteroseismology
To gain deeper insights into the population of the Gaia-observed $\beta$ Cep stars, we used the probabilistic identification of the dominant modes. In the following, we characterise the ensemble and subject all sample stars to one homogenous type of basic modelling by relying on a grid of stellar evolution models. The aim of this work is to understand the properties of the population of field $\beta$ Cep stars instead of doing detailed studies of only a few individual stars. Our approach is a more powerful way to understand global aspects of stellar evolution theory and pinpoint some of its possible shortcomings for the entire mass regime covered by the sample.
@@ 6.1 Distributions of the identified modes
The mean behaviour of the identified modes of the full sample over the wavelength is very similar to the hand-selected sample in Fig. 9. However, we find a slightly clearer separation between the two considered non-radial modes. Notably the distribution of $A_{T}/A_{\mathrm{BP}}$ for $l=1$ shown in Fig. 23 has a heavy tail in the TESS amplitude but not in the other photometric bands.
Figure 11: Amplitude ratio diagram based on $A_{\mathrm{RP}}/A_{\mathrm{BP}}$ and $A_{T}/A_{\mathrm{BP}}$ similar to Fig. 7. The colour-code shows the mode degree as determined with our probabilistic framework: $l=0$ (green), $l=1$ (blue), $l=2$ (orange). Grey symbols denote stars with $p<0.6$. The symbol size is proportional to the probability of the mode identification. Stars marked with black outlines constitute the underlying sample from Fig. 9.
@@
When going into details, we find a complicated pattern with several overlapping regions. Due to the three dimensionality of the distributions these overlap regions create fuzzy borders in the two-dimensional projection in Fig. 11. Stars to which our algorithm assigned low probabilities in the mode identification belong to the overlapping regions between the different degrees as also seen from the shown contours. An interesting feature is the group of high probability $l=1$ modes at the bottom of Fig. 11. This region creates the tail in the $A_{\mathrm{T}}/A_{\mathrm{BP}}$ amplitude ratios in Fig. 23. From the TESS data, we find that some of these stars have multiplets in which the Gaia observed frequency (and often also the main TESS frequency) is not the central frequency. Indeed, the multi-colour photometry also carries information on the intrinsic mode excitation amplitudes, aside from the mode geometry and inclination angle of the star. Our findings illustrate that the mode amplitudes are not equal among the prograde and retrograde components of the same $|m|$ within a multiplet for these heat-driven modes. This is in contrast to the case for stochastically excited modes in slowly-rotating low-mass stars, where the amplitudes allow for the derivation of the star’s inclination angle. Mode visibilities purely based on geometrical arguments are indeed not the only factor that determines the amplitudes (and hence detectability) within a multiplet for $\beta\,$Cep stars, as has been known since long from ground-based photometry.
When placing the stars with identified modes in the HRD, we find that the $\beta$ Cep stars with a dominant radial mode are mainly of lower mass ($\lesssim 15\,M_{\sun}$). The radial pulsators seem to concentrate in the later half of the main sequence in the overlap region between the $\beta$ Cep and SPB instability strips. In contrast, non-radial pulsators can be observed throughout the whole $\beta$ Cep mass range (cf. Fig. 22), although two stars with radial modes stand out due to their relatively high luminosity and mass.
@@ 6.2 Rotationally split multiplets and asymmetries
As introduced in Sect. 5.2, rotation splits non-radial modes into multiplets with splittings proportional to the averaged envelope rotation rate. We now return to the 14 stars with likely multiplets (not used in the definition set to construct the KDE in Sect. 5.3). We now have their probabilistic mode identification. Adding them to the 19 stars with deterministic identification from the full multiplet structure we had already from Sect. 5.2, they compose a sample of 33 stars with multiplets which we now use to study the stellar rotation properties. The full sample of these 33 stars with multiplets is shown in Figs. 42 and 44.
@@
In Fig. 12, we show the observed splittings against the estimated (projected) surface rotation rate. As expected, a positive correlation between the rotation rate and the splitting is observed. Yet, we recall that $\beta$ Cep stars can have strong radial differential rotation (Aerts et al., 2003; Pamyatnykh et al., 2004; Burssens et al., 2023) and the surface rotation rate may differ from the internal rotation rate at the fractional radius where the mode is most sensitive. Stars at $v\sin i=0$ km s-1 have undetectable line broadening in Gaia, yet we include them here as they are included as such in Gaia DR3.
@@
Figure 12: Observed mode splittings against the projected surface rotation rate ($f_{\mathrm{rot}}\sin i$, encircled in thin grey) or photometric surface rotation rate ($f_{\mathrm{rot}}$, encircled in thick solid black). The three lines show the splitting values when equal to the surface rotation rate (solid) as well as half (dotted) and twice (dashed) the surface rotation rate. The uncertainties on the splittings indicate the maximum difference of the individual splittings to the mean splitting of the multiplet and are a measure for the asymmetry. The colours show our deterministic identifications as $l=1$ (blue) and $l=2$ (orange) modes.
@@
We also mark the absolute asymmetries (maximum difference to the mean splitting) in Fig. 12. For most of the stars this asymmetry is small, though it is pronounced for a minority. A purely rotationally split multiplet would consist of evenly and symmetrically spaced components if the Coriolis force due to the mean envelope rotation has only a small influence. This allows for a first-order perturbative approach to compute the oscillation mode frequencies. However, second-order effects are relevant for $\beta$ Cep stars and include rotational deformation of the star (Saio, 1981). This effect shifts the split frequencies asymmetrically with respect to the central frequency (Dziembowski & Goode, 1992; Goupil et al., 2000). In addition, magnetic fields (e.g. Bugnet et al., 2021) or binarity (e.g. Guo, 2021) also influence the split multiplets, leading to asymmetries in the observed frequency spectra and enabling detailed asteroseismic inferences (Guo et al., 2024).
@@
Figure 13: Asymmetries of 33 rotationally split multiplets $\delta_{A}$ against their central frequency $f_{0}$ colour-coded by the projected surface rotation rate. Square symbols mark stars with measured photometric surface rotation rates ($f_{\mathrm{rot}}$). Symbols connected by a line belong to the same ($l=2$) multiplet. Black outlines indicate stars where the Gaia frequencies do not match the TESS frequencies within our threshold. The measurement uncertainties are within the symbol sizes.
@@
with $f_{0}$ the central frequency and $f_{\pm m}$ the frequencies of azimuthal order $m=\pm 1$. In Fig. 13, we show the 33 most obvious asymmetries occurring in the multiplets detected in our sample stars against the central frequency $f_{0}$ of the multiplet. Most of these splittings are positive as expected for rotational deformation (Guo et al., 2024). Negative asymmetries in $\beta$ Cep stars can hint towards internal magnetic fields (Mathis & Bugnet, 2023), especially for dipole p-modes. Our understanding of core magnetism in massive stars is still limited (Lecoanet et al. 2022; Vanlaer et al., in prep.), despite its potential impact on stellar evolution. A large sample of measured internal magnetic fields from asteroseismology would deliver constraints for evolutionary models, similarly to the recent advances in red giants (Li et al., 2023; Deheuvels et al., 2023).
@@ 6.3 Modelling the population of $\beta$ Cep stars
To understand the population of galactic $\beta$ Cep stars better and gain insights into the distributions of their physical properties with respect to the selected main pulsation mode, we applied grid modelling. In doing so, we used extra seismic information compared to classical grid modelling where one just relies on evolution tracks and the measured effective temperature and luminosity to determine stellar properties. Here, we used the $\beta$ Cep MESA/GYRE stellar evolution and pulsation grids presented by Burssens et al. (2023). The evolution grid was calculated for masses between 9 $M_{\sun}$ and 21.5 $M_{\sun}$ at solar metallicity with envelope mixing by internal gravity waves set to $D=1000\,$cm2 s-1 at the base of the envelope and increasing outwards as $\rho^{-1}$ and an exponentially decaying diffusive core overshoot with the parameter range $0.005\leq f_{\mathrm{ov}}\leq 0.035$. For details of the physical implementation, we refer the reader to Burssens et al. (2023).
@@
The required mode identification for the modelling is not limited to the degree $l$ but we also need the radial order $n$ and azimuthal order $m$. All models will be fit in the azimuthal order $m=0$ necessitating us to identify the central peak of split modes denoted here as $f_{0}$. For complete multiplets or radial pulsators, $f_{0}$ is measured. For the few quadrupole pulsators with incomplete multiplets, we assumed the highest peak of the central two peaks to be the one with $m=0$.
@@
The radial order can neither be deduced from the observations nor from a direct fit of the model. Rather, we fitted models with the radial orders $n\in\{0,1,2\}$ for radial pulsators and $n\in\{-2,-1,0,1,2\}$ for non-radial modes [^6] The f-mode ($n=0$) only exists for $l=2$ as the frequency goes to zero for dipole modes (e.g. Takata, 2012). using a $\chi^{2}$ minimisation over $\log T_{\mathrm{eff}}$, $\log L$, and $f_{0}$. In principle, the radial order can be larger but higher order modes have not been securely established in $\beta$ Cep stars, hence we did not probe them.
@@
Stars without a model within the $1\sigma$ uncertainty ellipse were excluded from the subsequent analysis, reducing the number of modelled stars to 119. For each star, we chose the best fitting model for each $n$ and compared the position of its other modes to the observed frequency spectra. Our focus laid on the relative positions of the main pulsation mode and other visible modes with respect to those predicted by the model. In most but not all cases our choice of $n$ coincides with the lowest $\chi^{2}$ value among the set of best fitting models. We show one illustrative example for the identification in Appendix C, including a detailed reasoning for our choice of $n$.
@@
We find the majority of the radial pulsators to pulsate in their fundamental mode (eight out of twelve modelled radial pulsators). Three radial pulsators pulsate in their first overtone and one stars is likely a second overtone pulsator. A similar picture emerges for the 107 non-radial pulsators. The dominant pulsation mode in 50 stars is the $p_{1}$-mode ($n=1$) and 24 pulsate in $g_{1}$ ($n=-1$). Only 29 stars pulsate in their second-order mode with 28 identified $p_{2}$ and one $g_{2}$ pulsators. The remaining four stars were identified to be dominant f-mode pulsators.
@@
For the final model selection and uncertainty estimation, we employed a MCMC approach with a $\chi^{2}$ merit function, similar to Fritzewski et al. (2024), to which we refer for details. Since we rely on a grid of stellar models with fixed input physics on some aspects (notably the level of envelope mixing, where the low-order modes have dominant probing power), we allow for an overall mismatch between observed and fitted frequencies (see Aerts et al., 2018, for an extensive discussion). In order to encapsulate this theoretical uncertainty caused by the fixed yet uncertain input physics adopted in the model grid, we follow the predictions for zonal p-modes by Aerts et al. (2018) in the $\beta\,$Cep mass regime and allow for a typical frequency uncertainty of 0.1 d-1. This captures any systematic errors in the predicted frequencies while allowing to recognise the correspondence between the observed and predicted frequency pattern (cf.  the procedure explained in in Appendix C).
Figure 14: Distribution of $\beta$ Cep parameters for the modelled stars. Each panel contains the histograms for the full sample (grey) and each degree $l=0$ (green), $l=1$ (blue), and $l=2$ (orange). From top to bottom, these panels show the initial mass distribution, the normalised core hydrogen fraction $X_{c}^{\prime}$, the stellar age in years, the relative convective core mass, the radius and the K-value.
@@
The histograms in Fig. 14 show distributions of the obtained masses, normalised core hydrogen fraction ($X_{c}^{\prime}$), age, relative convective core mass, radius, and horizontal versus radial displacement of the dominant mode. As expected from the distribution of $\beta$ Cep stars in the HRD in Fig. 3, the number of $\beta$ Cep stars declines with mass, with most $\beta$ Cep stars found in the range $10\leq M_{\mathrm{ini}}/M_{\sun}\leq 15$. Eight stars could not be modelled because they lie above the upper mass limit of $21.5\,M_{\sun}$, as shown by the evolutionary tracks in Fig. 3. In agreement with Fig. 22, radial pulsators are mostly found among the lower mass $\beta$ Cep stars.
The distribution of $X_{c}^{\prime}$ shows more structure when focusing on the pulsation degree. Overall, the majority of our sample is in their second half of the main-sequence evolution. In particular, the small group of radial pulsators are found in the second half of their main sequence evolution, while non-radial pulsators are found over the whole range.
@@
We do not show a distribution of the core overshoot because most of the stars are modelled with $f_{\mathrm{ov}}\approx 0.02$, which corresponds to the central value in the grid. Yet, the uncertainties span the whole range of $f_{\mathrm{ov}}$ covered by the grid, which can be explained by the limited sensitivity of the observed modes to the near-core region and the degeneracies between the mass, metallicity, and core overshoot when relying on just one mode (e.g., Ausseloos et al., 2004). The estimated core overshoot values result in a value for the mass-normalised convective mass ($M_{\mathrm{core}}/M_{\star}$). This inferred parameter is the more interesting quantity from a stellar evolution point of view and it is a function of both the luminosity and age. The majority of stars has a value in the range $0.2\leq M_{\mathrm{core}}/M_{\star}\leq 0.3$, in agreement with the distribution of initial mass shown in the top panel and the HRD. Notably the radial pulsators are concentrated around $M_{\mathrm{core}}/M_{\star}\approx 0.25$ due to their common evolutionary status. The few stars with a relatively high convective core mass belong to the higher-mass young O-stars well separated in the HRD (cf. Fig. 3). The radius distribution is peaked at intermediate radii, especially for dominant radial mode pulsators.
@@
TIC Gaia DR3 source ID Name $X_{c}^{\prime}$ $M_{ini}$ $\log t$ $M_{\mathrm{core}}/M$
@@
As a final distribution, we show in the bottom panel the distribution for the $K-$value, defined as $K\equiv GM/(2\pi f_{0})^{2}R^{3}$, for the non-radial modes. This dimensionless quantity is connected with the boundary conditions for the pulsation equations (Aerts et al., 2010, Chapter 3). In the case of non-radial modes, it represents the ratio between the horizontal and radial displacement of the fluid elements at the stellar surface caused by the mode (as this ratio is zero for radial modes, these are not included in the bottom panel of Fig. 14). This parameter has high values above unity (typically between 10 and 100) for high-order g-modes of observed SPB stars (De Cat & Aerts, 2002), as their modes dominantly cause horizontal displacements. Here, as expected for low-order modes, we find the opposite with dominantly radial displacements but with non-negligible contributions from horizontal motions. The values for $K$ range from about 0.03 to about 0.2 for the dominant mode of most of the Gaia $\beta$ Cep stars in our sample.
Figure 15: Normalised convective core mass $M_{\mathrm{core}}/M_{\mathrm{ini}}$ over the normalised core hydrogen mass fraction $X_{c}^{\prime}$. The stars are colour-coded by their primary degree: $l=0$ (green), $l=1$ (blue), and $l=2$ (orange). The shape represents the radial order: g2-modes are represented by thin diamonds, g1 by thin triangles, radial fundamental modes by triangles, p1 by circles and p2 by squares. The grey tracks in the background show the evolution for the indicated masses with $f_{\mathrm{ov}}=0.005$ (solid, labelled at low $X_{c}^{\prime}$) and $f_{\mathrm{ov}}=0.035$ (dashed, labelled at high $X_{c}^{\prime}$). The typical uncertainty is indicated in the upper right corner.
To unravel the correlation between the convective core mass and the core hydrogen fraction, we show these two parameters in Fig. 15. The overall distribution shows clearly the distinction between stars with $M_{\mathrm{core}}/M_{\star}\leq 0.3$ and these with more massive convective cores. The stars with higher relative core masses are typically more massive and younger as expected from stellar evolution theory. As seen from the evolutionary tracks of different masses, only the lowest mass $\beta$ Cep stars can be found among full breadth of the main sequence evolution, a consequence of the position of the instability strip lying closer to the zero-age main sequence for these stars.
In summary, we successfully modelled 119 $\beta$ Cep stars with a grid modelling approach adding the seismic information on the dominant mode to the two classical observables, effective temperature and luminosity. We provide distributions of their masses, ages, convective core masses, radii. As such, these distributions provide a first insight into the properties of the population of $\beta$ Cep stars. Of particular interest for future investigations on the mode excitation of this population is the connection between the stellar age and the degree of the dominant mode, if any. We give the best fitting model properties in Table 2. A discussion of the accuracy of the asteroseismic parameters in comparison with the input values from Gaia can be found in Appendix D.1.